# On the Relationship between Truth and Political Bias in Language Models

### v1

@@ Abstract
Language model alignment research often attempts to ensure that models are not only helpful and harmless, but also truthful and unbiased. However, optimizing these objectives simultaneously can obscure how improving one aspect might impact the others. In this work, we focus on analyzing the relationship between two concepts essential in both language model alignment and political science: truthfulness and political bias. We train reward models on various popular truthfulness datasets and subsequently evaluate their political bias. Our findings reveal that optimizing reward models for truthfulness on these datasets tends to result in a left-leaning political bias. We also find that existing open-source reward models (i.e. those trained on standard human preference datasets) already show a similar bias and that the bias is larger for larger models. These results raise important questions about both the datasets used to represent truthfulness and what language models capture about the relationship between truth and politics.
@@ 1 Introduction
Our results suggest that even training on supposedly objective datasets can lead to unforeseen bias.
@@ 2 Related Work
Prior work has extensively covered ways to ‘align’ models with human preferences (Bai et al., 2022; Casper et al., 2023), particularly the widely used technique of reinforcement learning from human feedback, or RLHF Stiennon et al. (2020). Other work has examined how truth is represented in language models (Burns et al., 2022; Azaria and Mitchell, 2023), sometimes in terms of embedding space geometry (Marks and Tegmark, 2023). The nature of truth, however, is philosophically complicated (Levinstein and Herrmann, 2024) and there are many open problems (Farquhar et al., 2023). Prior work has also found that LLMs have political biases (Motoki et al., 2023), and traced these biases’ connection to the political opinions in training data (Santurkar et al., 2023; Feng et al., 2023).
@@ 3 Experimental Setup
We use several datasets corresponding to different notions of factuality to train our reward models: TruthfulQA Lin et al. (2022), FEVER Thorne et al. (2018), SciQ Welbl et al. (2017), and a dataset we created of 4,000 basic LLM-generated facts and falsehoods about the world, using GPT-4 (OpenAI et al., 2023) and Gemini (Gemini Team et al., 2024). (See Appendix B for details of how we generated, validated and audited this last dataset.) To make the data suitable for reward modeling, which expects paired samples, we match a correct response to a query with an incorrect response for TruthfulQA, FEVER, and SciQ. For the generated dataset, we create random pairs of true and false statements. For datasets with multiple-choice options, we ensure that each question appears exclusively in either training or test.
@@
To test reward models for political bias, we use GPT-3.5-turbo (OpenAI, 2023) to generate TwinViews-13k, a dataset consisting of 13,855 pairs of left-leaning and right-leaning statements matched by topic. The model was instructed to keep the statements as similar as possible in style and length. We used generated statements because of the dearth of large topically matched datasets of political statement pairs; for example, the popular political compass test [^1] https://www.politicalcompass.org/test includes only a few statements. We extensively audited the generated statements to ensure their relevance and quality. Details of the prompt and the quality-assurance process, including a sample of the statement pairs (Table 4), can be found in Appendix A. We release the final TwinViews dataset publicly for use by the community.
Figure 2: “Truthful” reward models usually show a left-leaning political bias. The left three subplots show rewards assigned to TwinViews political statements by models fine-tuned on each truthfulness dataset, excluding explicitly political content found by our audit. Individual points show each run’s results, while the bar height shows the average. Note the presence of inverse scaling: Larger models usually skew further left. Results of Section 5.2’s n-gram experiment appear in the rightmost pane, showing no clear relationship to the neural models’ patterns.
@@
We clarify terminology with respect to the different model types here. A “base” model refers to a pre-trained LLM without any further fine-tuning, while a “vanilla” reward model is a base model fine-tuned on standard human preference datasets such as OpenAssistant Köpf et al. (2023), Anthropic Helpful-Harmless Bai et al. (2022), and OpenAI’s summarizing from human feedback data Stiennon et al. (2020). A “truthful” reward model is a base model fine-tuned on a truthfulness dataset.
For experiments on vanilla reward models, we evaluate RMs from RAFT222weqweasdas/hh-rlhf-rm-open-llama-3b (Dong et al., 2023), OpenAssistant [^3] OpenAssistant/reward-model-deberta-v3-large-v2 and UltraRM444openbmb/UltraRM-13b (Cui et al., 2023). For the truthful reward models, we train several RMs on each truthfulness dataset (Section 3) with weights initialized from the base 160M, 2.8B and 6.9B Pythia models (Biderman et al., 2023), conducting several runs on different splits ($80\%$ train, $20\%$ test) for robustness. (All runs are shown in Figure 2.) We also train a simple tri-gram baseline on each dataset for the analysis in Section 5.2 (see the rightmost pane of Figure 2). After training these models (details in Appendix E), we run inference on the TwinViews data to test whether the truthful reward models still show political bias.
@@ 4 Bias in Vanilla Reward Models
We first examine whether vanilla open-source reward models exhibit political bias. As discussed in Section 3, we evaluate with models from RAFT, OpenAssistant and UltraRM. We run inference with these models on the TwinViews statements and find that all models show a left-leaning political bias, as depicted in Figure 1. Notably, larger models also show greater bias, an example of inverse scaling (McKenzie et al., 2023). However, one caveat is that the datasets/training methods are different across these reward models. The results suggest that at least part of the left-leaning political bias observed in the literature Santurkar et al. (2023) could be due to biases introduced in reward-model training, which we believe is a new finding.
@@ 5 Bias in “Truthful” Reward Models
While vanilla reward models exhibit a clear political slant, these models are fine-tuned on datasets of subjective human preferences reflecting diverse goals Casper et al. (2023). Our objective is to minimize this subjectivity by training “truthful reward models”—reward models designed to give high scores to objectively truthful statements (e.g., basic everyday facts or scientific information) and low scores to false statements. As discussed in Section 3, we pursue this goal by fine-tuning various base Pythia models as reward models on each of the four truthfulness datasets, and evaluating the rewards they assign to the left and right TwinViews statements. Because any resulting political bias might be due to political content in the truthfulness datasets, we first systematically audit them for such content (in Section 5.1). We find very low rates of political content, but nevertheless exclude it from subsequent model training and analysis.
Training models on these cleaned datasets produces results shown in the left three panes of Figure 2. We found that our truthful reward models generally assign higher rewards to left-leaning statements than right-leaning ones (in 11 out of 12 cases). As with vanilla models, the degree of bias also usually increased with model size.
With fine-tuning datasets intended to be objective, these findings were unexpected. In Section 5.2, we use an n-gram baseline (shown in the rightmost pane of Figure 2) to consider another potential source of bias: stylistic features spuriously correlated with both truth status and political orientation. We find little support for this idea either, however, leaving the origin of the political bias shown in Figure 2 in need of further research.
@@ 5.1 Explicit Political Bias
Political content in truthfulness datasets may lead to political bias in models trained on them. However, our analysis shows that these datasets contain very little explicitly political content. We used two methods, building on a list of political topics from the Comparative Agendas Project Jones et al. (2019), to identify political content.
@@
First, we used a simple keyword matching approach. We generated potential political keywords with GPT-4, and used them to search for potential political content. We then manually labeled the flagged training examples. This method found that about $2\%$ of the data in TruthfulQA contains some political content, while less than $1\%$ of the data in the other datasets is politics-related. Specifically, SciQ includes 35 examples about climate change, and FEVER contains ten examples about politicians, though these are mostly factual.
As a robustness check, we also used GPT-3 to search for political content in a subset of 1000 examples from each dataset. [^5] We used GPT-3 because OpenAI’s API returns log-probabilities of arbitrary completions only for GPT-3 models. The results confirmed the low levels of explicitly political content. Details of both methods are given in Appendix D.
@@ 5.2 Stylistic Artifacts
Even after excluding explicitly political content, a left-leaning bias might arise from “stylistic” features of the truthfulness data. For instance, if negation words (e.g., “no,” “not”) are more prevalent in both false and right-leaning statements, the reward model might learn to associate these features, as with the length bias in some RMs Shen et al. (2023). We test this hypothesis with the n-gram baseline: If this simple model shows a political bias similar to that of the neural models, it would support the idea that those models’ bias stems from stylistic features of the datasets.
@@ 6 Bias Across Topics
Because both vanilla and “truthful” reward models show political bias, we used regression analysis to examine which topics or political issues exhibit the most bias. For both sets of models, we regressed the reward assigned to a TwinViews political statement on several predictors: the model, [^6] For the truthful models, each Pythia model fine-tuned on each dataset is a separate level of this variable, for 12 in total. the topic, the statement’s political lean, and the topic/political-lean interaction. All models are linear regression.
@@ 7 Conclusion
We investigated political biases in reward models, both vanilla open-source reward models and “truthful” reward models, and found a persistent left-leaning political bias across nearly all these models. This result is particularly surprising given the use of datasets designed to capture objective truth. Moreover, the size of the bias increases with model scale, in contrast to the usual pattern of improving capabilities. For the “truthful” models, we considered and attempted to rule out two explanations: explicit political content in truthfulness datasets and spurious relationships between truthfulness and stylistic features. Identifying the source of this bias is a promising direction for future research, and we hope these initial findings will encourage further investigation into the relationship between truthfulness and political bias in language models.
@@ 8 Limitations
Though the relationship between truth and political bias in language models is a timely and important topic, this study has certain limitations in addressing it. Firstly, datasets are an imperfect representation of truth and falsehood. Although there has been significant interest in identifying truthful directions in LLMs Marks and Tegmark (2023); Azaria and Mitchell (2023); Burns et al. (2022), recent work has shown that these findings are sensitive to simple perturbations, such as negation Farquhar et al. (2023); Levinstein and Herrmann (2024). Consequently, it is possible that the reward models are learning dataset artifacts rather than a true notion of truth versus falsehood. Nevertheless, it is valuable to understand how these artifacts may affect political bias. Secondly, our study focuses solely on reward models. While there are good reasons for this focus (reward models are a crucial component of the RLHF pipeline and their scalar outputs allow simple quantitative comparison of preferences), it still restricts what we can say about the rest of the alignment pipeline. Future research should explore how optimizing models through other alignment methods, such as direct preference optimization, or DPO (Rafailov et al., 2023), impacts the downstream model in more externally valid settings such as text generation.
@@ Appendix B Generated True/False Statements
statements. The statement should all be
@@
We generated several thousand of these statements in batches for maximum diversity. To minimize the number of duplicated statements, we then used k-means clustering on their sentence embeddings to find 2000 clusters and picked one true and one false statement from each cluster. Finally, we manually inspected several hundred samples to check whether the statements were properly true or false and apolitical. We show sample statements in Table 5.
@@ Appendix D Identifying Political Content
We first generated a list of keywords corresponding to various political topics via GPT-4. This list of keywords is included in supplementary material. Then, we manually went through any content matched by the keywords and classified whether it actually contained explicitly political content. We report the total number of topics tagged and removed in Table 3.
@@ Appendix E Model Training Details
We train all models on an NVIDIA A6000 GPU. All models are trained with an effective batch size of 128 and a learning rate of $4e{-5}$ for one epoch. The 2.8B and 6.9B parameter models are trained with PEFT, with hyperparameters $r=128$ and LoRA’s $\alpha=128$. All parameters of the 160M model were fine-tuned. We estimate each training run took between ten and thirty GPU minutes depending on the dataset size. With three model sizes, four datasets, and five iterations each, with an average of 20 minutes per run, we estimate our total computational budget was around 20 GPU hours.

### v2

@@ Abstract
Language model alignment research often attempts to ensure that models are not only helpful and harmless, but also truthful and unbiased. However, optimizing these objectives simultaneously can obscure how improving one aspect might impact the others. In this work, we focus on analyzing the relationship between two concepts essential in both language model alignment and political science: truthfulness and political bias. We train reward models on various popular truthfulness datasets and subsequently evaluate their political bias. Our findings reveal that optimizing reward models for truthfulness on these datasets tends to result in a left-leaning political bias. We also find that existing open-source reward models (i.e., those trained on standard human preference datasets) already show a similar bias and that the bias is larger for larger models. These results raise important questions about the datasets used to represent truthfulness, potential limitations of aligning models to be both truthful and politically unbiased, and what language models capture about the relationship between truth and politics.
@@ 1 Introduction
Our results suggest that even training on supposedly objective datasets can lead to unforeseen bias. We also release a dataset of 13,855 left-leaning and right-leaning partisan statements matched on topic for use by the community [^1] Code and data available here..
@@ 2 Related Work
We briefly cover three areas that our work relates to: AI alignment, LLM truthfulness, and political bias in LLMs.
2.1 Alignment
Prior work has extensively covered ways to ‘align’ models with human preferences (Bai et al., 2022; Casper et al., 2023), particularly the widely used technique of reinforcement learning from human feedback, or RLHF Stiennon et al. (2020). Recent methods like DPO (Rafailov et al., 2023) bypass creating an explicit reward model; however, alignment datasets may still contain biases depending on the annotators’ values and preferences (Kirk et al., 2024).
2.2 Truthfulness in LLMs
Other work has examined how truth is represented in language models (Burns et al., 2022; Azaria and Mitchell, 2023), sometimes in terms of embedding space geometry (Marks and Tegmark, 2023). The nature of truth, however, is philosophically complicated (Levinstein and Herrmann, 2024a). Several of these works present both theoretical and empirical challenges, leaving it an open question whether language models genuinely possess “truth representations" (Farquhar et al., 2023; Levinstein and Herrmann, 2024b). However, some approaches have shown promise in increasing truthfulness of LLMs by intervening on intermediate representations (Li et al., 2023; Chuang et al., 2024).
2.3 Political bias in LLMs
Prior work has also found that LLMs have political biases (Motoki et al., 2023; Bang et al., 2024), and traced these biases’ connection to the political opinions in training data (Santurkar et al., 2023; Feng et al., 2023). This literature generally finds a left-leaning bias in LLMs; however, there are some topics where LLMs respond with right-leaning perspectives (Perez et al., 2023). There have also been methods proposed to reduce the political bias of language models Liu et al. (2021).
Finally, there has been extensive research in political science on partisan differences in attitudes toward truth, such as misinformation (Baptista and Gradim, 2022) and trust in science (Cologna et al., 2024). Our work sits at the intersection of these areas of research, attempting to understand how truth and political views intersect with LLMs.
@@ 3 Experimental Setup
We use several datasets corresponding to different notions of factuality to train our reward models: TruthfulQA Lin et al. (2022), FEVER Thorne et al. (2018), SciQ Welbl et al. (2017), and a dataset we created of 4,000 basic LLM-generated facts and falsehoods about the world, using GPT-4 (OpenAI et al., 2023) and Gemini (Gemini Team et al., 2024). (See Appendix B for details regarding how we generated, validated and audited this last dataset.) FEVER is based on facts about entities extracted from Wikipedia. SciQ is based on scientific knowledge. TruthfulQA covers a variety of topics and was created with the goal of eliciting untruthful completions from LLMs. Finally, our generated data aimed to create the most obvious facts and falsehoods. Thus, our datasets span facts about entities (FEVER), scientific facts (SciQ), a diverse mix of difficult questions (TruthfulQA), and common sense facts (our generated data). To make the data suitable for reward modeling, which expects paired samples, we match a correct response to a query with an incorrect response for TruthfulQA, FEVER, and SciQ. For the generated dataset, we create random pairs of true and false statements. For datasets with multiple-choice options, we ensure that each question appears exclusively in either training or test.
@@
To test reward models for political bias, we use GPT-3.5 Turbo (OpenAI, 2023) to generate TwinViews-13k, a dataset consisting of 13,855 pairs of left-leaning and right-leaning statements matched by topic. The model was instructed to keep the statements as similar as possible in style and length. We used generated statements because of the dearth of large topically matched datasets of political statement pairs; for example, the popular political compass test [^2] https://www.politicalcompass.org/test includes only a few statements. We extensively audited the generated statements to ensure their relevance and quality. Details of the prompt and the quality-assurance process, including a sample of the statement pairs (Table 4), can be found in Appendix A. However, we note that using LLM generated data can lead to a variety of issues, such as the risk of agreement bias, and thus we would encourage users of this data to consider these limitations (see Section 8 for a more thorough discussion). We release the final TwinViews dataset publicly for use by the community.
Figure 2: “Truthful” reward models usually show a left-leaning political bias. The left three subplots show rewards assigned to TwinViews political statements by models fine-tuned on each truthfulness dataset, excluding explicitly political content found by our audit. We run five train/eval splits for each dataset and model. Individual points show results from each run, with blue points representing the average reward given to left-leaning statements and red points representing the average reward given to right-leaning statements. The red and blue bar heights show the average reward across all five runs (i.e. the average of the corresponding point values). Note the presence of inverse scaling: Larger models usually skew further left. Results of Section 5.2’s n-gram experiment appear in the rightmost pane, showing no clear relationship to the neural models’ patterns.
@@
Here we clarify terminology with respect to the different model types. A “base” model refers to a pre-trained LLM without any further fine-tuning, while a “vanilla” reward model is a base model fine-tuned (only) on standard human preference datasets such as OpenAssistant Köpf et al. (2023), Anthropic Helpful-Harmless Bai et al. (2022), and OpenAI’s summarizing from human feedback data Stiennon et al. (2020). A “truthful” reward model is a base model fine-tuned on a truthfulness dataset (with no preceding fine-tuning on human preference data).
For experiments on vanilla reward models, we evaluate RMs from RAFT333weqweasdas/hh-rlhf-rm-open-llama-3b (Dong et al., 2023), OpenAssistant [^4] OpenAssistant/reward-model-deberta-v3-large-v2 and UltraRM555openbmb/UltraRM-13b (Cui et al., 2023). These models were chosen due to their diversity in size and training data/methods, such that any measured political bias would be relatively generalizable. For the truthful reward models, we train several RMs on each truthfulness dataset (Section 3) with weights initialized from the base 160M, 2.8B and 6.9B Pythia models (Biderman et al., 2023), conducting several runs on different splits ($80\%$ train, $20\%$ test) for robustness. (All runs are shown in Figure 2.) We choose the Pythia models because their pretraining data is transparent and they cover a range of sizes, allowing us to understand how political bias scales with model size. We also train a simple tri-gram baseline on each dataset for the analysis in Section 5.2 (See the rightmost pane of Figure 2). After training these models (details in Appendix E), we run inference on the TwinViews data to test whether the truthful reward models still show political bias.
@@ 4 Bias in Vanilla Reward Models
We first examine whether vanilla open-source reward models exhibit political bias. As discussed in Section 3, we evaluate with reward models from RAFT, OpenAssistant and UltraRM. We run inference with these models on the TwinViews statements and find that all models show a left-leaning political bias, as depicted in Figure 1. Notably, larger models also show greater bias, an example of inverse scaling (McKenzie et al., 2023). However, one caveat is that the datasets/training methods are different across these reward models. The results suggest that at least part of the left-leaning political bias observed in the literature Santurkar et al. (2023) could be due to biases introduced in reward-model training, which we believe is a new finding.
@@ 5 Bias in “Truthful” Reward Models
While vanilla reward models exhibit a clear political slant, these models are fine-tuned on datasets of subjective human preferences reflecting diverse goals Casper et al. (2023). Our objective is to minimize this subjectivity by training “truthful reward models”—reward models designed to give high scores to objectively truthful statements (e.g., basic everyday facts or scientific information) and low scores to false statements. As discussed in Section 3, we pursue this goal by fine-tuning various base Pythia models as reward models on each of the four truthfulness datasets, and evaluating the rewards they assign to the left and right TwinViews statements. Because any resulting political bias might be due to political content in the truthfulness datasets, we first systematically audit them for such content (in Section 5.1). We find very low rates of political content, but nevertheless exclude it from subsequent model training and analysis. Training models on these cleaned datasets produces results shown in the left three panes of Figure 2. We found that our truthful reward models generally assign higher rewards to left-leaning statements than right-leaning ones (in 11 out of 12 cases). As with vanilla models, the degree of bias also usually increased with model size.
Given that fine-tuning datasets are intended to be objective, these findings were unexpected. In Section 5.2, we use an n-gram baseline (shown in the rightmost pane of Figure 2) to consider another potential source of bias: stylistic features spuriously correlated with both truth status and political orientation. We find little support for this idea either, however, leaving the origin of the political bias shown in Figure 2 in need of further research.
@@ 5.1 Explicit Political Bias
Political content in truthfulness datasets may lead to political bias in models trained on them. However, our analysis shows that these datasets contain very little explicitly political content. We used two methods, building on a list of political topics from the Comparative Agendas Project Jones et al. (2019) to identify political content.
@@
First, we used a simple keyword matching approach. We generated potential political keywords with GPT-4 and used them to search for potential political content. We then manually labeled the flagged training examples. This method found that about $2\%$ of the data in TruthfulQA contains some political content, while less than $1\%$ of the data in the other datasets is politics-related. Specifically, SciQ includes 35 examples about climate change, and FEVER contains 10 examples about politicians, though these are mostly factual.
As a robustness check, we also used GPT-3.5 to search for political content in a subset of 1000 examples from each dataset. [^6] We used GPT-3.5 because OpenAI’s API returns log-probabilities of arbitrary completions only for GPT-3.5 models. The results confirmed the low levels of explicitly political content. Details of both methods are given in Appendix D.
@@ 5.2 Stylistic Artifacts
Even after excluding content that is explicitly political, a left-leaning bias might arise from “stylistic” features of the truthfulness data. For instance, if negation words (e.g., “no,” “not”) are more prevalent in both false and right-leaning statements, the reward model might learn to associate these features, as with the length bias in some RMs Shen et al. (2023). We test this hypothesis with the n-gram baseline: If this simple model shows a political bias similar to that of the neural models, it would support the idea that those models’ bias stems from stylistic features of the datasets.
@@ 6 Bias Across Topics
Because both vanilla and “truthful” reward models show political bias, we used regression analysis to examine which topics or political issues exhibit the most bias. For both sets of models, we regressed the reward assigned to a TwinViews political statement on several predictors: the model, [^7] For the truthful models, each Pythia model fine-tuned on each dataset is a separate level of this variable, for 12 in total. the topic, the statement’s political lean, and the topic/political-lean interaction. All models are linear regressions.
@@ 7 Conclusion
We investigated political biases in reward models, both vanilla open-source reward models and “truthful” reward models, and found a persistent left-leaning political bias across nearly all these models. This result is particularly surprising given the use of datasets designed to capture objective truth. Moreover, the size of the bias increases with model scale, in contrast to the usual pattern of improving capabilities. For the “truthful” models, we considered and attempted to rule out two explanations: explicit political content in truthfulness datasets and spurious relationships between truthfulness and stylistic features. Identifying the source of this bias is a promising direction for future research, as well as understanding whether optimizing for truth leads to more or less political bias than other objectives.
More generally, this work connects to the increasing politicization of scientific facts, such as climate change (Hulme, 2009), and the problem of “truth decay" Kavanagh and Rich (2018) in the political sphere, which sit at the intersection of truth and politics. Finally, our results suggest a potential tension in achieving both truthful and unbiased models which has important implications for LLM alignment. We hope these initial findings will encourage further investigation into the relationship between truthfulness and political bias in language models.
@@ 8 Limitations
Our study has certain limitations, some inherent to notions of politics and truth, and some which we hope future work can investigate.
Politics is relative
It is difficult to create truly future-proof datasets of either apolitical factual statements or political statements, because what is considered political changes over time. Any seemingly factual issue may become politicized, as with climate change, or a political issue may cease to be controversial. In addition to being temporally localized, the definition of “political” content also varies between cultures, and our definitions come from a Western and especially US-centric perspective. We hope future work can audit truthfulness datasets for political content in a more expansive fashion. Adopting a broader notion of politics beyond the common left-right spectrum, would also help capture this rich context.
Difficulty of capturing truth
Datasets are an imperfect representation of truth and falsehood. Despite significant interest in identifying truthful directions in LLMs Marks and Tegmark (2023); Azaria and Mitchell (2023); Burns et al. (2022), recent work has found such directions sensitive to simple perturbations like negation Farquhar et al. (2023); Levinstein and Herrmann (2024a). It is thus possible that our reward models learn dataset artifacts rather than truth and falsehood as such. Nevertheless, it is valuable to understand how these artifacts affect political bias in practice. Similarly, depending on generated data to measure political bias also has limitations. Biases may be introduced from both the prompts used to generate the content and the LLMs themselves, stemming from datasets used and choices made in their pre- or post-training.
Only reward models
We study only reward models here. While there are good reasons for this focus (they are a crucial component of the RLHF pipeline and their scalar outputs allow simple quantitative comparison of preferences), it still restricts what we can say about other alignment methods. Future research should explore how methods like direct preference optimization, or DPO (Rafailov et al., 2023), impact models aligned with them.
@@ 9 Ethical Considerations
Acknowledgements
We would like to thank Yoon Kim for his helpful comments and feedback on this work.
@@ Appendix B Generated True/False Statements
statements. The statements should all be
@@
We generated several thousand of these statements in batches for maximum diversity. To minimize the number of duplicated statements, we then used k-means clustering on their sentence embeddings to find 2000 clusters and picked one true and one false statement from each cluster. Finally, we manually inspected several hundred samples to check whether the statements were properly true or false and apolitical, finding that between 90-95% are objectively true/false, with some statements being more subjective even though we prompted the model to only provide objective facts. We show sample statements in Table 5.
@@ Appendix D Identifying Political Content
We first generated a list of keywords corresponding to various political topics via GPT-4. This list of keywords is included in the supplementary material. Then, we manually went through any content matched by the keywords and classified whether it actually contained explicitly political content. We report the total number of topics tagged and removed in Table 3.
@@ Appendix E Model Training Details
We ran five train/test splits for each dataset and model to ensure robustness, with each split shuffling the order of the training examples. For the truthful datasets that came with prompts (SciQ and TruthfulQA), we simply used the questions provided as the prompts. For FEVER, since the topic was provided, we prompted the model with “Can you tell me a true statement about [TOPIC]?”, and for the generated true/false statements we prompted the model with “Can you tell me a true statement?”. This was to ensure consistency in that every dataset followed the Question-Answering format.
We train all models on an NVIDIA A6000 GPU. All models are trained with an effective batch size of 128 and a learning rate of $4e{-5}$ for one epoch. The 2.8B and 6.9B parameter models are trained with PEFT, with hyperparameters $r=128$ and LoRA’s $\alpha=128$. All parameters of the 160M model were fine-tuned. We estimate each training run took between 10 and 30 GPU minutes depending on the dataset size. With three model sizes, four datasets, and five iterations each, with an average of 20 minutes per run, we estimate our total computational budget was around 20 GPU hours.